OpenStudy (anonymous):

The sum of facing page numbers in a book is 943 . What are the page numbers? Show how you got answer?

OpenStudy (anonymous):

facing pages have consecutive page numbers let's denote one as x, so other is x+1 their sum is 943 x+(x+1)=943 2x+1=943 2x=942 x=942/2